simplify: square root sign 2 times square root sign 60
Answer by stanbon(75887)

simplify: square root sign 2 times square root sign 60
--------------------
[sqrt(2)] * [sqrt(60)] = sqrt(120) = sqrt(4*30) = 2sqrt(30)
==================
Cheers,
Stan H.